Circular No. 6672

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997co IN NGC 5125 The Beijing Astronomical Observatory (BAO) Supernova Survey (cf.

IAUC 6612, plus D.-w. Lu) reports the discovery of a supernova (mag ∼ 18.2) on an unfiltered CCD frame taken on May 29.7 UT with the BAO 0.6-m reflector, with a confirming image obtained under poor weather conditions on May 31. SN 1997co is located at α = 13^{h}24^{m}01^{s}.18, δ = +9^{o}42′22′′.9 (equinox 2000.0), which is 6′′.8 east and 14′′.5 south of the nucleus of NGC 5125. Unfiltered CCD images of the same field taken on May 16 (limiting mag ∼ 19.0) show no star at this position. S. Nakano, Sumoto, Japan, forwards the following position end figures and magnitude from M. Aoki, Tsukioka, Toyama (0.43-m f/6 reflector + CCD; limiting mag 18.5), obtained on May 31.502: 01^{s}.15, 23′′.2, V = 17.0.

L. Wang and J. C. Wheeler, University of Texas; and P. O. Shull, Oklahoma State University, write: “A low-resolution, low-signal-to-noise spectrogram of SN 1997co was obtained on May 31.2 UT with the 2.7-m telescope (+ LCS) at McDonald Observatory, showing this to be a type-II supernova with a conspicuous Hα P-Cyg profile. The expansion velocity measured at the minimum of the absorption trough is ∼ 5000 km/s.”

COMET C/1997 J1 (MUELLER) Extension to ephemeris on IAUC 6644 (elements MPC 29880):

1997TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
June 1 9^{h}26.^{m}80 +62^{o}40.^{′}3 2.611 2.326 62.^{o}6 22.^{o}8 12.8
6 9 21.51 +60 32.7 2.693 2.335 58.8 21.8 12.8
11 9 18.00 +58 33.2 2.774 2.345 55.0 20.8 12.9
16 9 15.81 +56 41.7 2.852 2.356 51.4 19.7 13.0
21 9 14.59 +54 57.7 2.928 2.369 47.8 18.5 13.1
26 9 14.11 +53 20.8 3.000 2.382 44.4 17.4 13.2
July 1 9 14.18 +51 50.5 3.068 2.397 41.1 16.2 13.2
6 9 14.69 +50 26.3 3.132 2.413 38.0 15.0 13.3
11 9 15.52 +49 07.8 3.190 2.431 35.2 13.9 13.4
16 9 16.59 +47 54.6 3.243 2.449 32.7 12.9 13.4
21 9 17.83 +46 46.2 3.290 2.469 30.5 12.1 13.5

Total visual magnitude estimates: May 13.02 UT, 13.2 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 25.88, 12.6 (K. Hornoch, Lelekovice, Czech Republic, 0.35-m reflector); 28.11, 12.5 (J. E. Bortle, Stormville, NY, 0.41-m reflector).

1997 May 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green